Two water pipes can fill a tank with water in 6 hours. The larger pipe working alone can fill the tank in 9 hours. How long will it take the smaller pipe working alone to fill the tank?
Answer by stanbon(75887)

Two water pipes can fill a tank with water in 6 hours. The larger pipe working alone can fill the tank in 9 hours. How long will it take the smaller pipe working alone to fill the tank?
---------
Together DATA:
time = 6 hrs/job ; rate = 1/6 job/hr
--------------
Larger pipe DATA:
time = 9 hrs/job ; rate = 1/9 job/hr
--------------
Smaller pipe DATA:
time = x hrs/job ; rate = 1/x job/hr
---------------
Equation:
rate + rate = together rate
1/x + 1/9 = 1/6
Multiply thru by 54x to get:
54 + 6x = 9x
3x = 54
x = 18 hrs (time for the smaller pipe to do the job)
======================================================
Cheers,
Stan H.